# Mean field equation and Green’s function on the Tate curve

### 16 Mar 2026 v1

@@ 3 Mean field equation on the finite quotient
$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ (3.1)
@@
$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ (3.2)
@@
$=p^{-d}\sum_{\begin{subarray}{c}0\leq k\leq m-1\\ 1\leq a_{0}\leq p-1\\ 0\leq a_{i}\leq p-1\\ 1\leq i\leq d-1\end{subarray}}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x))$
@@
$\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\rho p^{d}\boldsymbol{e}_{y}$ (3.4)
@@ 3.1 Structure of the solution
In this part, we present structural properties of the solution to (3.4). These properties are observed from numerical results.
@@
If $\boldsymbol{u}$ solves (3.4), then $u_{i}\leq d\text{ln}p,\forall i$. The equality holds if and only if $p=2,m=1$.
@@
$\sum_{i}e^{u_{i}}=p^{d}$
then $e^{u_{i}}\leq p^{d}$, thus $u_{i}\leq d\text{ln}p$ for all $i$. And the equality holds if and only if $p=2,m=1$ as there is one term in the sum. ∎
@@
$u_{y}$ is the only smallest in $\boldsymbol{u}$.
@@
Let $m=\text{min}_{i}\ u_{i}$. Suppose $\exists\ i\neq y$, s.t. $u_{i}=m$. From (3.5) with $i\neq y$,
$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=0$
@@
Let $\boldsymbol{u}=\begin{pmatrix}\boldsymbol{u}_{1}\\ \boldsymbol{u}_{2}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, $\boldsymbol{\sigma}=\begin{pmatrix}&&&1\\ 1&&&\\ &\ddots&&\\ &&1&\end{pmatrix}\in\mathbb{R}^{(p-1)p^{d-1}\times(p-1)p^{d-1}}$ be the shift matrix on blocks of size $(p-1)p^{d-1}$, and define $\sigma_{k}(\boldsymbol{u})=\begin{pmatrix}\boldsymbol{u}_{1}\\ \vdots\\ \boldsymbol{\sigma}\boldsymbol{u}_{k}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}$. When $i\neq j$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ as $\boldsymbol{B}_{ij}=p^{-|i-j|}\boldsymbol{1}\boldsymbol{1}^{\top}$ when $i\neq j$.
@@
$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{k}}=\boldsymbol{b}_{k}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
@@
$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{\sigma}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho\boldsymbol{\sigma}e^{\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
@@
$\begin{cases}\displaystyle\sum_{j\neq k}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\boldsymbol{A}_{1k}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\boldsymbol{A}_{kk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{\sigma}\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k}\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\boldsymbol{A}_{mk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
Thus $\sigma_{k}(\boldsymbol{u})$ is a solution of $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\sigma_{k}(\boldsymbol{b})$. ∎
Lemma 3.4 and Lemma 3.5 imply certain symmetric properties: given one solution, we can construct many solutions by folding or rotating the solution. However, if the solution is radial symmetric, then Lemma 3.4 and Lemma 3.5 are trivial. From numerical results we can see that there indeed have radial symmetric solutions, and by the following theorem we proved that there only exist radial symmetric solutions when $\rho$ is small.
@@
When $\rho\leq p^{-d}-p^{-d-m}$, denote $u_{y}(x)$ be the solution of (3.4), then $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$. Moreover, if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ and $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$, which means that $u(x)$ is radial symmetric.
@@
$=$ $\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
$-(\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j})u(x)$
@@
$\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}$ (3.7)
$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d}\sum_{i=0}^{d-2}p^{i}$
@@
$=$ $p^{2d-1}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1}$
which coincides with the negative of the diagonal entries of $\boldsymbol{B}$. Note that
@@
$=$ $\sum_{\begin{subarray}{c}a_{0}\neq b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}p^{2}u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\neq b_{d-1}\end{subarray}}p^{2(d-1)}u(z)$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-1)u(z)$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}=b_{1}\\ a_{2}\neq b_{2}\\ a_{3}\dots a_{d-1}\end{subarray}}(p^{4}-p^{2})u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2})u(z)$
@@
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2(d-2)})u(z)$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$
@@
$p^{d}Du(x)=$ $\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)$
$+\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$
$-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)$
@@
$\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$ (3.8)
$+\rho p^{d}e^{u(x)}-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)=0$
Denote $u(x)=u_{l}(b_{0},b_{1},\dots,b_{d-1})$, by fixing $b_{0},b_{1},\dots,b_{d-2}$ and varying $b_{d-1}$, we see that $u(x)$ satisfies (LABEL:1_4). In addition, the function $f(x)=\rho p^{d}e^{x}-kx,\ x\leq d\text{ln}p$ is monotonic when $p^{2d}\rho\leq k$. Then by Lemma 3.2, if $p^{2d}\rho\leq p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-1})$ is constant for all $b_{d-1}$ except at $x=y$. Hence for $x\neq y$, we denote $u(x)$ by $u_{l}(b_{0},b_{1},\dots,b_{d-2})$.
When $v_{p}(x)=l\neq v_{p}(y)$, (LABEL:1_4) becomes
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{v_{p}(z)=l}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$
$+p(p^{2(d-1)}-p^{2(d-2)})u_{l}(b_{0},b_{1},\dots,b_{d-2})+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}$
$-(p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0$

### 08 Apr 2026 v2

@@ 3 Mean field equation on the finite quotient
$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}.$ (3.1)
@@
$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p}).$ (3.2)
@@
$=p^{-d}\sum_{\begin{subarray}{c}0\leq k\leq m-1\\ 1\leq a_{0}\leq p-1\\ 0\leq a_{i}\leq p-1\\ 1\leq i\leq d-1\end{subarray}}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x)).$
@@
$\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\rho p^{d}\boldsymbol{e}_{y}.$ (3.4)
@@ 3.1 Structure of the solution
If $\boldsymbol{u}$ solves (3.4), then $u_{i}\leq d\ln p,\forall i$. The equality holds if and only if $p=2,m=1$.
@@
$\sum_{i}e^{u_{i}}=p^{d},$
then $e^{u_{i}}\leq p^{d}$, thus $u_{i}\leq d\ln p$ for all $i$. And the equality holds if and only if $p=2,m=1$ as there only has one term in the sum. ∎
@@
$u_{y}$ is the only smallest one in $\boldsymbol{u}$.
@@
Let $m=\min_{i}u_{i}$. Suppose $\exists\ i\neq y$, s.t. $u_{i}=m$. From (3.5) with $i\neq y$,
$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=0,$
@@
Let $\boldsymbol{u}=\begin{pmatrix}\boldsymbol{u}_{1}\\ \boldsymbol{u}_{2}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$ where $\boldsymbol{u}_{i}\in\mathbb{R}^{(p-1)p^{d-1}}$, $\boldsymbol{\sigma}=\begin{pmatrix}&&&1\\ 1&&&\\ &\ddots&&\\ &&1&\end{pmatrix}\in\mathbb{R}^{(p-1)p^{d-1}\times(p-1)p^{d-1}}$ be the shift matrix on blocks of size $(p-1)p^{d-1}$, and define $\sigma_{k}(\boldsymbol{u})=\begin{pmatrix}\boldsymbol{u}_{1}\\ \vdots\\ \boldsymbol{\sigma}\boldsymbol{u}_{k}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}$ as $\boldsymbol{A}_{ij}$ is circulant. Furthermore, when $i\neq j$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ as $\boldsymbol{B}_{ij}=p^{-|i-j|}\boldsymbol{1}\boldsymbol{1}^{\top}$ when $i\neq j$. Therefore, we have
$\boldsymbol{\sigma}\boldsymbol{A}_{ii}=\boldsymbol{A}_{ii}\boldsymbol{\sigma}$ (3.6)
and
$\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ (3.7)
when $i\neq j$.
Remark 3.5.
[…]
@@
$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{k}}=\boldsymbol{b}_{k},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}.\end{cases}$
@@
$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{\sigma}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho\boldsymbol{\sigma}e^{\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m},\end{cases}$
@@
$\begin{cases}\displaystyle\sum_{j\neq k}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\boldsymbol{A}_{1k}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1},\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\boldsymbol{A}_{kk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{\sigma}\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k},\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\boldsymbol{A}_{mk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}.\end{cases}$
by (3.6) and (3.7). Thus $\sigma_{k}(\boldsymbol{u})$ is a solution of $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\sigma_{k}(\boldsymbol{b})$. ∎
Lemma 3.4 and Lemma 3.6 imply certain symmetric properties: since $\boldsymbol{b}=p^{d}\boldsymbol{e}_{y}$, so given one solution, we can construct many solutions by folding or rotating the solution unless the solution is radially symmetric. And by the following theorem we proved that there only exist radially symmetric solutions when $\rho$ is small.
@@
When $\rho\leq p^{-d}-p^{-d-m}$, denote $u_{y}(x)$ be the solution of (3.4), then $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, which means that the solution is radially symmetric. Moreover, if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ and $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$.
@@
$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
$-\left(\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}\right)u(x).$
@@
$\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}$ (3.9)
$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d-1-j}\sum_{j=1}^{d-1}p^{2j}$
$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d}\sum_{j=0}^{d-2}p^{j}$
@@
$=$ $p^{2d-1}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1},$
which coincides with the negative of the diagonal entries of $\boldsymbol{B}$.
Note that
@@
$=$ $\sum_{\begin{subarray}{c}a_{0}\neq b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}p^{2}u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}p^{2(d-1)}u(z)$
$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-1)u(z)$
$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}=b_{1}\\ a_{2}\neq b_{2}\\ a_{3}\dots a_{d-1}\end{subarray}}(p^{4}-p^{2})u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2})u(z)$
@@
$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2(d-2)})u(z)$
$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z),$
@@
$p^{d}Du(x)=$ $\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)$
$+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x),$
@@
$\rho p^{d}e^{u(x)}+\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)$ (3.10)
$+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)=0.$
Denoting $u(x)=u_{l}(b_{0},b_{1},\dots,b_{d-1})$, by fixing $b_{0},b_{1},\dots,b_{d-2}$ and varying $b_{d-1}$, we see that $u(x)$ satisfies (LABEL:eq:equation_for_x_neq_y). In addition, the function $f(x)=\rho p^{d}e^{x}-kx,\ x\leq d\ln p$ is monotone when $p^{2d}\rho\leq k$. Then by Lemma 3.2, if $p^{2d}\rho\leq p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-1})$ is constant for all $b_{d-1}$ except at $x=y$. Hence for $x\neq y$, we denote $u(x)$ by $u_{l}(b_{0},b_{1},\dots,b_{d-2})$.
When $v_{p}(x)=l\neq v_{p}(y)$, (LABEL:eq:equation_for_x_neq_y) becomes
$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=l\end{subarray}}u(z)$
$+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})+p(p^{2(d-1)}-p^{2(d-2)})u_{l}(b_{0},b_{1},\dots,b_{d-2})$
$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0.$